OpenStudy (anonymous):

2 square root x-3 + 4 = x+1

OpenStudy (anonymous):

\[2\sqrt{x-3} + 4 = x+1\]

OpenStudy (anonymous):

always true!

OpenStudy (slaaibak):

2sqrt(x-3)=x-3 square both sides: 4x - 12 = x^2 - 6x + 9 x^2 - 10x + 21 = 0 (x-7)(x-3) = 0 x=7 or x=3

OpenStudy (anonymous):

best check your answers